OpenStudy (anonymous):

The manufacturing cost of a cell phone was $124. Sellers first marked up the cost by 20 percent. Sales were poor, so sellers decreased the markup to 15 percent. Calculate the old and new selling prices.

OpenStudy (anonymous):

\[old .M.P.=\frac{ 120 }{100 }\times 124=148.80 $\] \[New M.P.=\frac{ 100-15 }{ 100 }\times 148.80=?\]